### v2

@@ 6 Empirical applications
In this section we provide two empirical applications, one for measuring GE effect and the second one for externally valid site selection.
6.1 An implementation guide for measuring GE effects
Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Allcott et al., 2025). [^8] Across all AEA articles in 2015-2025, about $10\%$ of the experimental papers combine experimental results with structural models (i.e., one-third of the papers that use observational in combination with experimental estimates). Because large-scale experimentation are often infeasible due to cost constraints (Muralidharan and Niehaus, 2017) these models are meant to extrapolate effects from small-scale experiments to learn general-equilibrium (GE) effects. This extrapolation typically relies on other experiments or observational data.
We present a (step-by-step) illustration below when these models are calibrated using previous studies conducted in a different country.
Question
Consider a researcher evaluating a cash transfers for sending children to school in rural Kenya where such a program is not in place. Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects. Researchers have access to preliminary estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), which we may be concerned, lack external validity.
6.1.1 Step 1: model description
The first step for a researcher is to describe the modeling choices. We present a stylized model of school choice from Bonhomme and Weidner (2022) and Todd and Wolpin (2006).
@@ 6.1.1 Step 1: model description
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (13)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, let
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
@@
$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (14)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$. Under simple market clearing for labor supply and demand, we can show [^9] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (15)
where $d$ denotes the slope of the demand curve divided by total hours of work.
6.1.2 Step 2: parametrization and transparency on biases
Given the model, the following step for researchers is to define the parameters of interest. Here, we define the parameters of interest as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (16)
@@ 6.1.2 Step 2: parametrization and transparency on biases
where the explicit form of $w_{0}(\theta)$ follows from (15).
We treat $d$ as known for simplicity, using meta-analyses on demand elasiticities, and $y_{0}=1.5$ because computed experimentally in Kenya by Egger et al. (2022). [^10] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000). [^11] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when available.
In this example, the primary concern is bias in the observational estimator $\tilde{\theta}^{\mathrm{obs}}$ constructed as described below, abstracting from potential biases in $y_{0},d$ which, for simplicity, we assume to be second-order. When these biases are considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$. This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which biases’ concerns drive the design; these considerations can be collected in a pre-analysis plan.
6.1.3 Step 3: observational study estimates
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, and we will be focusing here on the effects on female students.
@@ 6.1.3 Step 3: observational study estimates
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
@@
We observe heterogeneity in both the sensitivity weights $\omega$ the precision of the observational estimates.
6.1.4 Step 4: describing the candidate experiments
The following step is to list the feasible experiments, subject to budget constraints.
@@ 6.1.4 Step 4: describing the candidate experiments
In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three possible experiments for a given sample size $n$:
@@
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
@@
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{1,2\}$: Two-arm design. The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$. They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
We consider two scenarios. In the first, the researcher (and oracle) can choose one or both experiments. In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$). Regret is defined accordingly, conditional on the set of feasible designs.
For simplicity, we calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$. In practice, researchers may use pilot studies to calibrate the experimental variances when available.
@@
Experiment type Identified parameter Choice Variables Calibrated standard error Description
@@
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
@@
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design. Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
@@
6.1.5 Step 5: experimental design
Figure 4: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are feasible. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
Figure 5: Optimal shrinkage weights and design choice as functions of the total sample size $n$. Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence. Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT. Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs. observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
Figure 6: Worst–case regret by design. Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization. In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks. Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias. Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
The final step is to design the experiment and report it in a pre-analysis plan.
Choice of the experiment when only one treatment arm is available
Figure 5 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses $n\approx 500$, CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Choice of the estimator
In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest. Figure 5 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
Sample size allocation when both arms are feasible
Figure 4 reports the sample allocation implied by the regret-optimal design when both treatment arms are feasible. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Regret comparisons
Figure 6 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only for small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
Implications
When only one arm is feasible, CCT is preferable only once the total sample size is sufficiently large; otherwise UCT is preferred. This pattern reflects the interaction between bias and variance: for small samples, CCT’s bias advantage is not large enough to compensate for its higher variance relative to UCT, whereas at larger samples the bias reduction becomes dominant. Existing designs based on variance considerations would fail to provide solutions that balance bias and variance comparisons.
When feasible, it is instead optimal to run two arms with highly unbalanced sample, reflecting the relative contribution of the variance for each treatment arm.
Following these steps, researchers can similarly report their design in a pre-analysis plan.
6.2 Measuring the performance of our method for site selection

### v3

@@ 6 Empirical applications
6.1 Experimental design for General Equilibrium effects
Because large-scale experimentation is often infeasible due to cost constraints (Muralidharan and Niehaus, 2017), researchers often combine experimental and observational data to extrapolate effects from small-scale experiments to learn general-equilibrium effects (e.g. Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022). The goal of this subsection is to present a step-by-step practitioner’s guide with an example, and compare our regret-optimal design to a standard variance-minimizing (Neyman) benchmark.
Consider a researcher evaluating conditional cash transfers for sending children to school in rural Kenya. Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects. Researchers have access to preliminary (external) estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), although these preliminary estimates may lack external validity in Kenya.
Step 1: model description
The first step for a researcher is to describe how observational/external and experimental estimates in Kenya are jointly used for estimation. Here, we consider a stylized model of school choice following Bonhomme and Weidner (2022), Todd and Wolpin (2006).
@@ 6.1 Experimental design for General Equilibrium effects
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{2}-\xi_{3}-\xi_{1})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (24)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{2}-\xi_{3}-\xi_{1})$ is without loss and simplifies expressions below. Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{3}W-\xi_{1}Y-\xi_{2}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
Model for GE effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages which we write as functions of a common transfer $t$ to all eligible households in the population. Specifically for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, for a common transfer $t$, denote the potential individual income and wage [^13] Because we focus on the effect of a common transfer, we keep a scalar notation for $t$ for expositional convenience. Alternatively, here $y(t),w(t)$ can also be interpreted as exposure mappings (Aronow et al., 2017) and $t$ can be defined as a $n$ dimensional vector of transfers.
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\hskip 18.49988ptW_{i}(t)=w(t)+\varepsilon_{Wi}.$
@@
$\mathbb{E}\!\left[\left.\frac{\partial\,\Pr(S{=}1\mid Y(t),W(t),t)}{\partial t}\right|_{t=0}\right]=\;\phi_{0}\cdot\big(\,\xi_{2}+\xi_{1}y_{0}-\xi_{3}w_{0}\,\big),$ (25)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{1}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$. Under standard market clearing for labor supply and demand, we write [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{2}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{3}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{2}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{3}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{2}}{\phi_{0}\,\xi_{3}-d},$ (26)
where $d$ denotes the slope of the demand curve divided by total hours of work. Wages in equilibrium depend both on the demand’s slope and the supply effect, captured through $\phi_{0}\xi_{2}$ (assuming that all non-students are employed for a constant number of hours).
Step 2: parametrization and transparency on biases
The following step is to define the parameters of interest, here defined as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{1},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{3},$ (27)
@@
where the explicit form of $w_{0}(\theta)$ follows from (26), a smooth function of the parameters.
We treat $d$ as known for simplicity, using meta-analyses on demand elasticities, and $y_{0}=1.5$ based on experimental estimates in Kenya from Egger et al. (2022). [^15] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000). [^16] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when this data is available.
Here, the primary concern is bias in $\tilde{\theta}^{\mathrm{obs}}$, abstracting from potential biases in $y_{0},d$. When biases of $(y_{0},d)$ are also considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$, which is allowed in our framework. This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which bias concerns drive the design; these considerations can be collected in a pre-analysis plan.
Step 3: observational study estimates
Using experimental data from the large-scale Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ through probit with standard controls (age, distance to school, eligibility, year, highest grade achieved) and use these estimates to obtain $\tilde{\theta}^{\mathrm{obs}}$. Estimation is conducted separately by gender; we will focus here on the effects on female students.
@@
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sqrt{\Sigma_{jj}^{\mathrm{obs}}}$
@@
Step 4: candidate experiments
@@
In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three stylized experiments or combinations thereof:
@@
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies the average marginal effect of income on school attendance, $\theta_{1}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial Y}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right].$ The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to PROGRESA.
@@
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies the average marginal effect of the stipend on school attendance, $\theta_{2}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial t}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right],$ with variance $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{3\}$: Youth job program (wage shock). Last, consider a job program for teenagers, that hires teenagers with small randomized differential wages. The program identifies $\theta_{3}$, holding constant $\theta_{1}$ and $\theta_{2}$
The per-unit variance in the experiment is calibrated to the corresponding entry of $\sigma^{2}$ in observational study, appropriately multiplied by the observational study sample size.
We consider two scenarios. First, the researcher may run only one of the three programs. Second, the researcher may run combinations of two of the three programs allocating a fixed total sample size to each, but not all three simultaneously, reflecting higher fixed costs and institutional constraints. For instance, the job program requires a different implementation infrastructure than a cash-transfer program; at the same time, conducting conditional and unconditional cash transfers in the same experimental location is often infeasible due to fairness concerns (e.g. Wingfield et al., 2023), which raises costs for each cash-transfer arm.
Remark 7 (Additional constraints).
In some contexts additional constraints may apply. For instance, we may be able to only run either $\{1,2\}$ or $\{3\}$, or only run either $\{1\}$ or $\{2\}$ but not $\{3\}$. Researchers can directly include these constraints in our framework. ∎
@@
Experiment type Identified parameters Choice Variables Calibrated standard errors Description
@@
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $\frac{6.56e^{-5}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
@@
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $\frac{9.86e^{-4}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{3\}$ $\theta_{3}$ $\gamma_{3}$ $\frac{1.01e^{-3}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Job program (Wage shock). Teenager jobs program with (small) randomized wages differentials.
$\{j,j^{\prime}\}$ $(\theta_{j},\theta_{j^{\prime}})$ $\gamma_{j},\gamma_{j^{\prime}},n_{j}$ $\frac{\sqrt{\Sigma_{jj}}}{\sqrt{n_{j}/n_{\mathrm{obs}}}}$, $\frac{\sqrt{\Sigma_{j^{\prime}j^{\prime}}}}{\sqrt{(n_{\mathrm{tot}}-n_{j})/n_{\mathrm{obs}}}}$ Combination of two of the three possible experiments. Choose two treatment arms and allocate optimally sample size, with given total experimental sample size $n_{\mathrm{tot}}$.
@@
Step 5: experimental design
Figure 3: Optimal experimental sample sizes $n_{j}$ allocated to each arm (UCT, CCT, and Job program) as a function of the total experimental sample size $n_{\text{tot}}$. The left panel considers designs with only one treatment arm, while the right panel allows two treatment arms.
Figure 4: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$. Left panel: one feasible experiment; right panel: two feasible experiments.
Finally, we design the experiment, with total number of experimental observations $n_{\mathrm{tot}}$; let $n_{\mathrm{obs}}$ denote instead the number of units in the original Mexican PROGRESA experiment.
Choice of the experiment and sample size
In Figure 3 we show the optimal sample size allocation as a function of the total experimental sample size, under the constraints that either one (left panel) or two (right panel) treatment arms can be selected. When the total sample size is small and only one arm can be chosen, the optimal design selects the UCT arm. Intuitively, the UCT strikes a good balance between sensitivity and precision of the experimental estimate when $n_{\text{tot}}$ is small.
When two treatment arms can be selected, the optimal design initially chooses both UCT and CCT for $n_{\mathrm{tot}}\leq 500$. In this region, almost all of the experimental sample is allocated to the CCT arm (which is more sensitive but noisier), with a smaller fraction allocated to the more precise UCT arm. For total sample sizes above five-hundred, the optimal design instead selects the job program as the main experiment, reflecting the high sensitivity of the estimand to the wage effect (Appendix Figure 12 reports the bias and variance regret of the design). With two experiments and large sample sizes, roughly 90% of the total sample is allocated to the job program and the remaining fraction to the UCT arm, using the precise UCT effect to help balance variance against bias.
In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest. Appendix Figure 13 collects the estimated shrinkage weight across designs. We find that the shrinkage weight is above 0.7, and closer to one for most cases. This shows how the experimental variation gets significantly higher weight.
Main bias and variance regret comparisons
In Figure 4 we compare the proposed optimal design to the variance-optimal (Neyman) design in terms of both variance and bias. This comparison clarifies what is gained by our procedure relative to a benchmark that chooses the experiment, shrinkage weights, and sample size solely to minimize the variance of the estimated effect $\hat{\tau}$ (and which therefore achieves the smallest possible variance). We find that the variance ratio between the Neyman and optimal designs (the inverse of the variance regret) is close to one for all sample sizes and converges to one as the sample size grows. For experimental sample $n_{\mathrm{tot}}=1000$, the variance of the Neyman and our proposed procedure is only about $14\%$ smaller when choosing one treatment arm and $35\%$ when choosing two treatment arms. In contrast, the bias of the Neyman procedure is about 1.4 times larger for one treatment arm and about 4.5 times larger with two. As a result, the adaptation regret of the Neyman allocation is similarly 1.4 and 4.5 times larger than our proposed design. These comparisons illustrate the benefit of our method: its variance is only slightly larger than the Neyman variance, while the bias is more than four times smaller for two treatment arms.
6.2 Site selection for external validity